OpenStudy (anonymous):

If x varies directly as y, and x = 7 1/2 when y = 10, find x when y = 4. direct variation?

OpenStudy (lgbasallote):

x:y = 15/2: 10 = x: 4 10x = 30 x = 3

OpenStudy (anonymous):

y = kx ← constant variation 4 = 8k k = 4/8 k = ½ Finally. y = x/2 y = ½ [If x = 1] I hope this helps!

OpenStudy (apoorvk):

becca its the same as the circle diameter question.. see hafiz's solution above.!